I will be 12 years older than my brother. I will be 28 years younger than my dad. If you add all our ages together, the sum will equal 154. How old will i be on my birthday?
Answer by lwsshak3(11628)

I will be 12 years older than my brother. I will be 28 years younger than my dad. If you add all our ages together, the sum will equal 154. How old will i be on my birthday?
***
let x=dad's age
x-28=son's age
x-28+12
..
x+x-28+x-28+12=154
3x-44=154
3x=198
x=66
x-28=38
How old will i be on my birthday? 38